# Qingdao 266237, China

### v1

@@ 5 MPSU from Symmetric-Key Techniques
In this section, we introduce a new primitive called multi-party secret-shared random oblivious transfer (mss-ROT), then we utilize its two-choice-bit version to construct a SK-MPSU based on oblivious transfer and symmetric-key operations in the standard semi-honest model.
@@ 5.1 Multi-Party Secret-Shared Random Oblivious Transfer
The abstraction of the general version of mss-ROT is inspired by the generation of Beaver Triples in multi-party setting. In the two-party GMW protocol, a Beaver Triple is produced by two role-switching executions of 1-bit ROT [ALSZ13]. As a result, $P_{1}$ holds $a_{1},b_{1},c_{1}$ and $P_{2}$ holds $a_{2},b_{2},c_{2}$, such that $(a_{1}\oplus a_{2})\cdot(b_{1}\oplus b_{2})=c_{1}\oplus c_{2}$. In the multi-party setting, the functionality of Beaver Triple is extended to let each party $P_{i}$ hold $a_{i},b_{i},c_{i}$ with the correlation $(\bigoplus_{i=1}^{m}a_{i})\cdot(\bigoplus_{i=1}^{m}b_{i})=\bigoplus_{i=1}^{m}c_{i}$. Given that $(\bigoplus_{i=1}^{m}a_{i})\cdot(\bigoplus_{i=1}^{m}b_{i})$ can be written as $(2-m)(\bigoplus_{i=1}^{m}a_{i}\cdot b_{i})+\bigoplus_{1\leq i\leq j\leq m}(a_{i}+a_{j})\cdot(b_{i}+b_{j})$, multi-party Beaver Triple reduces to pairwise two-party Beaver Triples. In consideration of the simple fact that the functionality of ROT can be interpreted as $r_{0}\oplus r_{b}=b\cdot\Delta$, where $\Delta=r_{0}\oplus r_{1}$, we follow the approach to producing multi-party Beaver Triples to extend ROT into multi-party setting, which is to let each party $P_{i}$ to individually input a choice bit $b_{i}$ and receive the output shares $r_{i}$ and $\Delta_{i}$ such that $r_{1}\oplus\cdots\oplus r_{m}=(b_{1}\oplus\cdots\oplus b_{m})\cdot(\Delta_{1}\oplus\cdots\oplus\Delta_{m})$.
In the context of MPSU, we only need two-choice-bit version of mss-ROT, which allows a more efficient construction with less pairwise invocations of two-party ROT compared to the general version. To elaborate, the two-choice-bit version of mss-ROT (for simplicity, we call it mss-ROT for short hereafter) allows two parties $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ to hold their choice bits $b_{0}$ and $b_{1}$ s.t. $r_{1}\oplus\cdots\oplus r_{m}=(b_{0}\oplus b_{1})\cdot(\Delta_{1}\oplus\cdots\oplus\Delta_{m})$. We give the formal functionality in Figure 10 and the detailed construction in Figure 11.
@@
It is easy to see that our construction essentially boils down to performing ROT pairwise. As one of the benefits, we can utilize the derandomization technique [Bea91] to bring most tasks forward to the offline phase. And the correctness and security of the mss-ROT protocol stems from the correctness and security of ROT. For the complete proof, refer to Appendix 0.E.1.
@@ 5.2 Construction of Our SK-MPSU
We now turn our attention to construct a SK-MPSU. The construction of our SK-MPSU follows the high-level ideas we introduced in the technical overview and is formally presented in Figure 12.
@@
Protocol $\Pi_{\mathsf{SK\text{-}MPSU}}$ securely implements $\mathcal{F}_{\mathsf{mpsu}}$ against any semi-honest adversary corrupting $t<m$ parties in the $(\mathcal{F}_{\mathsf{bssPMT}},\mathcal{F}_{\mathsf{mss\text{-}rot}},\mathcal{F}_{\mathsf{ms}})$-hybrid model, where $P_{1}=P_{\textnormal{ld}}$.
The security proof of Theorem 5.2 is deferred to Appendix 0.E.2.
  • Complexity Analysis. We provide a comprehensive complexity analysis for our SK-MPSU protocol in Appendix 0.F.2. The total costs are summarized below:
    • Online phase. The online computation complexity per party is $O(m^{2}n)$. The online communication complexity of $P_{1}$ is $O(\gamma mn+(l+\kappa)m^{2}n)$. The online communication complexity of $P_{j}$ is $O(\gamma mn+m^{2}n+(l+\kappa)mn)$. The online round complexity is $O(\log\gamma+m)$.
  • Parameters. $m$ parties $P_{1},\cdots,P_{m}$, where $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ provide inputs as shares of the choice bit, $\text{ch}_{0},\text{ch}_{1}\in[m]$. The message length $l$.
  • Functionality. On input $b_{0}\in\{0,1\}$ from $P_{\text{ch}_{0}}$ and $b_{1}\in\{0,1\}$ from $P_{\text{ch}_{1}}$,
    • Sample $r_{2},\cdots r_{m},\Delta_{1},\Delta_{2},\cdots,\Delta_{m}\leftarrow\{0,1\}^{l}$ and give $(r_{j},\Delta_{j})$ to $P_{j}$ for $2\leq j\leq m$.
    • If $b_{0}\oplus b_{1}=0$, compute $r_{1}=\bigoplus_{j=2}^{m}r_{j}$, else compute $r_{1}=\Delta_{1}\oplus(\bigoplus_{j=2}^{m}(r_{j}\oplus\Delta_{j}))$. Give $r_{1}$ to $P_{1}$.
@@
  • Parameters. $m$ parties $P_{1},\cdots,P_{m}$. where $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ provide inputs as shares of the choice bit, $\text{ch}_{0},\text{ch}_{1}\in[m]$. We use $J$ to denote the set of indices for the remaining parties who do not provide a choice-bit input, i.e., $J=[m]\setminus\{\text{ch}_{0},\text{ch}_{1}\}$. The message length $l$.
  • Inputs. $P_{\text{ch}_{0}}$ has input $b_{0}\in\{0,1\}$ and $P_{\text{ch}_{1}}$ has input $b_{1}\in\{0,1\}$.
@@
      $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ invoke $\mathcal{F}_{\mathsf{rot}}$ twice: First, $P_{\text{ch}_{0}}$ acts as receiver with input $b_{0}$ and $P_{\text{ch}_{1}}$ acts as sender. $P_{\text{ch}_{0}}$ receives $r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{b_{0}}\in\{0,1\}^{l}$. $P_{\text{ch}_{1}}$ receives $r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{0},r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{1}\in\{0,1\}^{l}$; Second, $P_{\text{ch}_{1}}$ acts as receiver with input $b_{1}$ and $P_{\text{ch}_{0}}$ acts as sender. $P_{\text{ch}_{1}}$ receives $r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{b_{1}}\in\{0,1\}^{l}$. $P_{\text{ch}_{0}}$ receives $r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{0},r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{1}\in\{0,1\}^{l}$.
@@
      For $j\in J$: $P_{\text{ch}_{0}}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{\text{ch}_{0}}$ acts as receiver with input $b_{0}$ and $P_{j}$ as sender without input. $P_{\text{ch}_{0}}$ receives $r_{j,{\text{ch}_{0}}}^{b_{0}}\in\{0,1\}^{l}$. $P_{j}$ receives $r_{j,{\text{ch}_{0}}}^{0},r_{j,{\text{ch}_{0}}}^{1}\in\{0,1\}^{l}$.
@@
      For $j\in J$: $P_{\text{ch}_{1}}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{\text{ch}_{1}}$ acts as receiver with input $b_{1}$ and $P_{j}$ as sender. $P_{\text{ch}_{1}}$ receives $r_{j,{\text{ch}_{1}}}^{b_{1}}\in\{0,1\}^{l}$. $P_{j}$ receives $r_{j,{\text{ch}_{1}}}^{0},r_{j,{\text{ch}_{1}}}^{1}\in\{0,1\}^{l}$.
@@
      For $j\in J$: $P_{j}$ samples $\Delta_{j}\leftarrow\{0,1\}^{l}$ and computes $r_{j}=r_{j,{\text{ch}_{0}}}^{0}\oplus r_{j,{\text{ch}_{1}}}^{0}$. $P_{j}$ sends $u_{j,{\text{ch}_{0}}}=\Delta_{j}\oplus r_{j,{\text{ch}_{0}}}^{0}\oplus r_{j,{\text{ch}_{0}}}^{1}$ to $P_{\text{ch}_{0}}$, and $u_{j,{\text{ch}_{1}}}=\Delta_{j}\oplus r_{j,{\text{ch}_{1}}}^{0}\oplus r_{j,{\text{ch}_{1}}}^{1}$ to $P_{\text{ch}_{1}}$, then outputs $(r_{j},\Delta_{j})$.
@@
      $P_{\text{ch}_{0}}$ computes $\Delta_{\text{ch}_{0}}=r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{0}\oplus r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{1}$ and $r_{\text{ch}_{0}}=\bigoplus_{j\in J}(r_{j,{\text{ch}_{0}}}^{b_{0}}\oplus b_{0}\cdot u_{j,{\text{ch}_{0}}})\oplus r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{0}\oplus r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{b_{0}}\oplus b_{0}\cdot\Delta_{\text{ch}_{0}}$ ($\cdot$ denotes bitwise-AND between the repetition code of $b_{0}$ and $u_{j,{\text{ch}_{0}}}$, which are both strings of length $l$. Similarly hereinafter), then outputs $(r_{\text{ch}_{0}},\Delta_{\text{ch}_{0}})$. $P_{\text{ch}_{1}}$ computes $\Delta_{\text{ch}_{1}}=r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{0}\oplus r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{1}$ and $r_{\text{ch}_{1}}=\bigoplus_{j\in J}(r_{j,{\text{ch}_{1}}}^{b_{1}}\oplus b_{1}\cdot u_{j,{\text{ch}_{1}}})\oplus r_{{\text{ch}_{1}},{\text{ch}_{0}}}^{0}\oplus r_{{\text{ch}_{0}},{\text{ch}_{1}}}^{b_{1}}\oplus b_{1}\cdot\Delta_{\text{ch}_{1}}$, then outputs $(r_{\text{ch}_{1}},\Delta_{\text{ch}_{1}})$.
@@
  • Parameters. $m$ parties $P_{1},\cdots,P_{m}$. Size $n$ of input sets. The bit length $l$ of set elements. Cuckoo hashing parameters: hash functions $h_{1},h_{2},h_{3}$ and number of bins $B$. An OKVS scheme $(\mathsf{Encode_{H}},\mathsf{Decode_{H}})$. A collision-resisitant hash function $\mathsf{H}(x):\{0,1\}^{l}\to\{0,1\}^{\kappa}$.
@@
      Hashing to bin. For $1\leq j\leq m$, $P_{j}$ does $\mathcal{C}_{j}^{1},\cdots,\mathcal{C}_{j}^{B}\leftarrow\mathsf{Cuckoo}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$ and $\mathcal{T}_{j}^{1},\cdots,\mathcal{T}_{j}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$.
@@
      Multi-party secret-shared random oblivious transfers. For $1\leq i<j\leq m,1\leq b\leq B$: $P_{1},\cdots,P_{j}$ invoke $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ where $P_{i}$ acts as $P_{\text{ch}_{0}}$ with input $e_{i,j}^{b}$ and $P_{j}$ acts as $P_{\text{ch}_{1}}$ with input $e_{j,i}^{b}$. For $1\leq d<j$, $P_{d}$ receives $r_{d,ji}^{b},\Delta_{d,ji}^{b}\in\{0,1\}^{l+\kappa}$ and computes $u_{d,j}^{b}=\bigoplus_{i=1}^{j-1}r_{d,ji}^{b}$. $P_{j}$ receives $r_{j,ji}^{b},\Delta_{j,ji}^{b}\in\{0,1\}^{l+\kappa}$ and computes $u_{j,j}^{b}=\bigoplus_{i=1}^{j-1}r_{j,ji}^{b}\oplus(\mathsf{Elem}(\mathcal{C}_{j}^{b})\|\mathsf{H}(\mathsf{Elem}(\mathcal{C}_{j}^{b})))$ if $\mathcal{C}_{j}^{b}$ is not corresponding to an empty bin, otherwise chooses $u_{j,j}^{b}$ at random. $\mathsf{Elem}(\mathcal{C}_{j}^{b})$ denotes the element in $\mathcal{C}_{j}^{b}$.

### v4

@@ 5 MPSU from Symmetric-Key Techniques
    In this section, we introduce a new primitive called multi-party secret-shared random oblivious transfer (mss-ROT), then we utilize it to build SK-MPSU based on oblivious transfer and symmetric-key operations in the standard semi-honest model.
@@ 5.1 Multi-Party Secret-Shared Random Oblivious Transfer
    In Figure 5, the ROT functionality can be interpreted as $r_{0}\oplus r_{b}=b\cdot\Delta$, where $\Delta=r_{0}\oplus r_{1}$. As the aforementioned two-choice-bit ROT, it can be consider to secret-share the choice bit $b$ between two parties and interpreted as $r_{0}\oplus r_{b}=(b_{0}\oplus b_{1})\cdot\Delta$. We further extend this secret-sharing idea to multiple parties, allowing not only any non-empty subset of the involved parties to secret-share $b$, but also any non-empty subset of the involved parties to secret-share $\Delta$.
    To build our SK-MPSU, we only need two-choice-bit version of mss-ROT. For simplicity, we will refer to this as mss-ROT hereafter. The formal functionality [^10] We choose to let the parties input the shares of $\Delta$ instead of outputting them, as the former functionality simply implies the latter. is given in Figure 9 and the detailed construction is given in Figure 10.
@@
    It is easy to see that our construction essentially boils down to performing ROT pairwise. As one of the benefits, we can utilize the derandomization technique [Bea91] to bring most tasks forward to the offline phase. And the correctness and security of the mss-ROT protocol stems from the correctness and security of ROT. For the complete proof, refer to Appendix 0.D.1.
@@ 5.2 Construction of Our SK-MPSU
    We now turn our attention to construct a SK-MPSU. The construction follows the high-level ideas we introduced in the technical overview and is formally presented in Figure 11.
@@
    Protocol $\Pi_{\mathsf{SK\text{-}MPSU}}$ securely implements $\mathcal{F}_{\mathsf{mpsu}}$ against any semi-honest adversary corrupting $t<m$ parties in the $(\mathcal{F}_{\mathsf{bssPMT}},\mathcal{F}_{\mathsf{mss\text{-}rot}},\mathcal{F}_{\mathsf{ms}})$-hybrid model.
    The proof of Theorem 5.2 is in Appendix 0.D.2. We provide a comprehensive complexity analysis for our SK-MPSU / LG and an exhaustive comparison between them in Appendix 0.E.
    • Parameters. $m$ parties $P_{1},\cdots,P_{m}$, where $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ provide inputs as shares of the choice bit, $\text{ch}_{0},\text{ch}_{1}\in[m]$. We use $J=\{j_{1},\cdots,j_{d}\}$ to denote the set of indices for the parties who provide inputs as shares of $\Delta$, where $d\leq m$. The message length $l$.
    • Functionality. On input $b_{0}\in\{0,1\}$ from $P_{\text{ch}_{0}}$, $b_{1}\in\{0,1\}$ from $P_{\text{ch}_{1}}$, and $\Delta_{j}$ from each $P_{j}$ where $j\in J$,
      • Sample $r_{i}\leftarrow\{0,1\}^{l}$ and give $r_{i}$ to $P_{i}$ for $2\leq i\leq m$.
      • If $b_{0}\oplus b_{1}=0$, compute $r_{1}=\bigoplus_{i=2}^{m}r_{i}$, else compute $r_{1}=\bigoplus_{i=2}^{m}r_{i}\oplus(\bigoplus_{j\in J}\Delta_{j})$. Give $r_{1}$ to $P_{1}$.
@@
    • Parameters. $m$ parties $P_{1},\cdots,P_{m}$. where $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ provide inputs as shares of the choice bit, $\text{ch}_{0},\text{ch}_{1}\in[m]$. We use $J=\{j_{1},\cdots,j_{d}\}$ to denote the set of indices for the parties who provide inputs as shares of $\Delta$, where $d\leq m$. The message length $l$.
    • Inputs. $P_{\text{ch}_{0}}$ has input $b_{0}\in\{0,1\}$ and $P_{\text{ch}_{1}}$ has input $b_{1}\in\{0,1\}$. Each $P_{j}$ for $j\in J$ has input $\Delta_{j}$,
@@
        For $1\leq i\leq m$, $P_{i}$ initializes $r_{i}=0$. If $ch_{j=0,1}\in J$, $P_{ch_{j}}$ sets $r_{ch_{j}}=b_{j}\cdot\Delta_{ch_{j}}$ ($\cdot$ denotes bitwise-AND).
@@
        For $j\in J$: $P_{\text{ch}_{0}}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{\text{ch}_{0}}$ acts as receiver with input $b_{0}$ and $P_{j}$ as sender. $P_{\text{ch}_{0}}$ receives $u_{j,{\text{ch}_{0}}}^{b_{0}}\in\{0,1\}^{l}$. $P_{j}$ receives $u_{j,{\text{ch}_{0}}}^{0},u_{j,{\text{ch}_{0}}}^{1}\in\{0,1\}^{l}$. $P_{j}$ updates $r_{j}=r_{j}\oplus u_{j,{\text{ch}_{0}}}^{0}$ and sends $\Delta^{\prime}_{j,{\text{ch}_{0}}}=u_{j,{\text{ch}_{0}}}^{0}\oplus u_{j,{\text{ch}_{0}}}^{1}\oplus\Delta_{j}$ to $P_{\text{ch}_{0}}$. $P_{\text{ch}_{0}}$ updates $r_{\text{ch}_{0}}=r_{\text{ch}_{0}}\oplus u_{j,{\text{ch}_{0}}}^{b_{0}}\oplus b_{0}\cdot\Delta^{\prime}_{j,{\text{ch}_{0}}}$.
@@
        For $j\in J$: $P_{\text{ch}_{1}}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{\text{ch}_{1}}$ acts as receiver with input $b_{1}$ and $P_{j}$ as sender. $P_{\text{ch}_{1}}$ receives $u_{j,{\text{ch}_{1}}}^{b_{1}}\in\{0,1\}^{l}$. $P_{j}$ receives $u_{j,{\text{ch}_{1}}}^{0},u_{j,{\text{ch}_{1}}}^{1}\in\{0,1\}^{l}$. $P_{j}$ updates $r_{j}=r_{j}\oplus u_{j,{\text{ch}_{1}}}^{0}$ and sends $\Delta^{\prime}_{j,{\text{ch}_{1}}}=u_{j,{\text{ch}_{1}}}^{0}\oplus u_{j,{\text{ch}_{1}}}^{1}\oplus\Delta_{j}$ to $P_{\text{ch}_{1}}$. $P_{\text{ch}_{1}}$ updates $r_{\text{ch}_{1}}=r_{\text{ch}_{1}}\oplus u_{j,{\text{ch}_{1}}}^{b_{1}}\oplus b_{1}\cdot\Delta^{\prime}_{j,{\text{ch}_{1}}}$.
@@
        If $\lvert I=\{ch_{0},ch_{1}\}\cup J\rvert<m$, each $P_{i}$ ($i\in I$) samples $r^{\prime}_{i,i^{\prime}}$ and sends $r^{\prime}_{i,i^{\prime}}$ to each $P_{i^{\prime}}$ ($i^{\prime}\in[m]\setminus I$) and updates $r_{i}=r_{i}\oplus r^{\prime}_{i,i^{\prime}}$. $P_{i^{\prime}}$ updates $r_{i^{\prime}}=r_{i^{\prime}}\oplus r^{\prime}_{i,i^{\prime}}$.
@@
        For $1\leq i\leq m$, $P_{i}$ outputs the final value of $r_{i}$.
@@
    • Parameters. $m$ parties $P_{1},\cdots,P_{m}$. Size $n$ of input sets. The bit length $l$ of set elements. Cuckoo hashing parameters: hash functions $h_{1},h_{2},h_{3}$ and number of bins $B$. A collision-resisitant hash function $\mathsf{H}(x):\{0,1\}^{l}\to\{0,1\}^{\kappa}$.
    • Inputs. Each party $P_{i}$ inputs $X_{i}=\{x_{i}^{1},\cdots,x_{i}^{n}\}\subseteq\{0,1\}^{l}$.
@@
        Hashing to bins. $P_{1}$ does $\mathcal{T}_{1}^{1},\cdots,\mathcal{T}_{1}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{1})$. For $1<j\leq m$, $P_{j}$ does $\mathcal{C}_{j}^{1},\cdots,\mathcal{C}_{j}^{B}\leftarrow\mathsf{Cuckoo}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$ and $\mathcal{T}_{j}^{1},\cdots,\mathcal{T}_{j}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$.
@@
        Multi-party secret-shared random oblivious transfers. For $1\leq i<j\leq m,1\leq b\leq B$: $P_{min(2,i)},\cdots,P_{j}$ invoke $\mathcal{F}_{\mathsf{mss\text{-}rot}}$ where $P_{i}$ acts as $P_{\text{ch}_{0}}$ with input $e_{i,j}^{b}$ and $P_{j}$ acts as $P_{\text{ch}_{1}}$ with input $e_{j,i}^{b}$. For $1<j^{\prime}\leq j$, $P_{j^{\prime}}$ samples $\Delta_{j^{\prime},ji}^{b}\leftarrow\{0,1\}^{l+\kappa}$ and inputs $\Delta_{j^{\prime},ji}^{b}$ as shares of $\Delta$. For $min(2,i)\leq d<j$, $P_{d}$ receives $r_{d,ji}^{b}\in\{0,1\}^{l+\kappa}$ and computes $u_{d,j}^{b}=\bigoplus_{i=1}^{j-1}r_{d,ji}^{b}$. $P_{j}$ receives $r_{j,ji}^{b}\in\{0,1\}^{l+\kappa}$ and computes $u_{j,j}^{b}=\bigoplus_{i=1}^{j-1}r_{j,ji}^{b}\oplus(\mathsf{Elem}(\mathcal{C}_{j}^{b})\|\mathsf{H}(\mathsf{Elem}(\mathcal{C}_{j}^{b})))$ if $\mathcal{C}_{j}^{b}$ is not corresponding to an empty bin, otherwise chooses $u_{j,j}^{b}$ at random. $\mathsf{Elem}(\mathcal{C}_{j}^{b})$ denotes the element in $\mathcal{C}_{j}^{b}$.